# Hierarchical autoregressive neural networks in three-dimensional statistical system

### v1

@@ 4.3 HAN versus Wolff algorithm: timings
It is also instructive to compare the speed of our HAN 3D algorithm with that of the Wolff cluster algorithm. Here we are not aiming for a precise comparison, since we used in-house code (with some optimizations) for the Wolff algorithm. We shall make this comparison just to give the reader a rough idea of a HAN timing. The exact comparison of the algorithms is difficult, also due to their very different nature. For example, Wolff algorithm in principle is sequential; however, one can still run multiple Markov Chains simultaneously on the GPU. Both, the Wolff algorithm and HAN, were run on the same device: a high-performance Nvidia GeForce 4090 GPU.
@@
Summarizing this rough comparison, the same precision for the observable was obtained with the cluster algorithm $\sim 70$ times faster than with our HAN 3D model (training and generation). It is worth noting that running the Wolff algorithm on a single CPU core is a few hundred times slower than using a GPU in our case. This is because of the small autocorrelation time, which allows for massive parallelization speed-up.
@@ 4.4 Efficiency of training in 2D and 3D
In Figure 7, we plot the ESS as a function of training epochs for those two versions of the HAN algorithm. We used the same hyperparameters for dense networks (2 layers) and the same learning rate 0.005. For each system, we run the training 5 times (curves of the same color). One can observe that ESS grows faster for 3D HAN than in 2D. The timings are the following: 500 epochs of training the 3D HAN last around 8.4 seconds and around 10 seconds for the 2D HAN.
@@ 5 Summary
In this paper, we have presented the hierarchical autoregressive network (HAN) algorithm suited for 3-dimensional spin systems and compared it with three other architectures: conventional variational autoregressive networks (VAN), autoregressive convolutional networks (PixelCNN), and a gated version of autoregressive convolutional networks (gated PixelCNN). We have performed several tests concerning the efficiency of 3D algorithms using classical Ising model: i) scaling with the system size, ii) comparison with the Wolff algorithm (concerning the evaluation of physical observables as well as the timings). We have found that HAN 3D is the fastest of all the architectures tested and allows good training quality (measured $ESS\approx 0.95$) in the system $8\times 8\times 8$. A rough comparison shows that for such a system, the Wolff cluster algorithm is much faster (factor $\sim 70$) on the same device. On the other hand, Wolff algorithm (and any other MCMC method) does not allow for estimation of thermodynamical observables like free energy and entropy. We calculated those using HAN (combined with NIS).
In our opinion, the fact that autoregressive networks can be used to calculate observables inaccessible using standard MCMC algorithms should be the main reason for the development of them. This manuscript is a step towards developing methods to calculate entropy, quantum entanglement entropy, and classical mutual information in 3D systems. Those quantities have been challenging to obtain using other methods, like tensor networks or Monte Carlo based on the Jarzynski equality [17]. The main obstacle for using NIS in 3D (and 4D) is the scalability of the neural samplers with system size. Although HAN is a good first step in this direction, much improvement is needed. Such improvement may come with new architectures developed for general-purpose Machine Learning.
@@ Acknowledgments
Computer time allocation ’plgnglft’ on the Athena supercomputer hosted by AGH Cyfronet in Kraków, Poland was used through the Polish PLGRID consortium. T.S. and D.Z. kindly acknowledge support of the Polish National Science Center (NCN) Grant No. 2021/43/D/ST2/03375. P.K. acknowledges support of the Polish National Science Center (NCN) grant No. 2022/46/E/ST2/00346. D.Z. acknowledges the support of the Research Support Module under the program Excellence Initiative - Research University at the Jagiellonian University. This research was partially funded by the Priority Research Area Digiworld under the program Excellence Initiative – Research University at the Jagiellonian University in Kraków.

### v2

@@ 4.3 HAN versus Wolff algorithm: timings
It is instructive to compare the speed of our HAN 3D algorithm with that of the Wolff cluster algorithm. Here, we are not aiming for a precise comparison, since we used the in-house code (with some optimizations) for the Wolff algorithm. We shall make this comparison just to give the Reader a rough idea of HAN timings. The exact comparison of algorithms is difficult, also due to their very different nature. For example, the Wolff algorithm, in principle, is sequential; however, one can still run multiple Markov Chains simultaneously on the GPU. Both the Wolff algorithm and HAN were run on the same device: a high-performance Nvidia GeForce 4090 GPU.
@@
In summary, the same precision for the observable was obtained with the cluster algorithm $\sim 70$ times faster than with our HAN 3D model (training and generation). It is worth noting that running the Wolff algorithm on a single CPU core is a few hundred times slower than using a GPU in our case. We could use this ”embarrassingly parallel” implementation because of a very small (compared to the number of collected samples) thermalization time, i.e. time needed to reach statistical equilibrium. Were this time longer, it would add an overhead that could substantially reduce the advantage of using parallelization on a slower machine. Also, the size of the systems that can be simulated in this way is constrained by the memory of the GPU.
@@ 4.4 Efficiency of training in 2D and 3D
In Figure 8, we plot the ESS as a function of training epochs for those two versions of the HAN algorithm. We used the same hyperparameters for dense networks (2 layers) and the same learning rate of 0.005. For each system, we run the training 5 times (curves of the same color). One can observe that ESS grows faster for 3D HAN than in 2D. The timings are the following: 500 epochs of training the 3D HAN last around 8.4 seconds, and around 10 seconds for the 2D HAN.
@@ 5 Summary
In this paper, we have presented the hierarchical autoregressive network (HAN) algorithm suited for 3-dimensional spin systems and compared it with three other architectures: dense variational autoregressive networks (VAN), autoregressive convolutional networks (PixelCNN), and a gated version of autoregressive convolutional networks (gated PixelCNN). The latter was never used in the context of spin generation. Although the 3D HAN algorithm uses the same idea as its 2D version, the code required significant changes concerning data manipulation when generating the spin configuration.
We have performed several tests concerning the efficiency of 3D algorithms using the classical Ising model: i) scaling of generation time with the system size, ii) comparison with the Wolff algorithm (evaluation of physical observables as well as the timings). We have found that HAN 3D is the fastest of all the neural architectures tested and allows good training quality (measured $ESS\approx 0.95$) for the system $8\times 8\times 8$. A rough comparison shows that for such a system, the Wolff cluster algorithm is much faster (factor $\sim 70$) on the same device. Using HAN (or other architectures considered in this work), it is straightforward to obtain thermodynamic observables such as free energy or entropy. We compared our results with the Wang-Landau algorithm, finding good agreement.
In our opinion, the fact that autoregressive networks can be used to calculate observables often hardly accessible using standard MCMC algorithms, like mutual information or entanglement entropy, should be the main reason for their development. This manuscript is a step towards developing methods to calculate such observables in 3D systems. Those quantities have been challenging to obtain using other methods, such as tensor networks or Monte Carlo, based on the Jarzynski equality [25]. The main obstacle to using NIS in 3D (and 4D) is the scalability of the neural samplers with system size. Although HAN is a good first step in this direction, much improvement is needed. Such an improvement may come from new architectures developed for general-purpose Machine Learning.
@@ Acknowledgments
We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016395. T.S. and D.Z. kindly acknowledge the support of the Polish National Science Center (NCN) Grant No. 2021/43/D/ST2/03375. P.K. acknowledges the support of the Polish National Science Center (NCN) grant No. 2022/46/E/ST2/00346. The study was funded by ”Research support module” as part of the ”Excellence Initiative – Research University” program at the Jagiellonian University in Kraków. This research was partially funded by the Priority Research Area Digiworld under the program Excellence Initiative – Research University at the Jagiellonian University in Kraków.
Appendix A Implementation details
In the Listing 1, we show the snippet of the main function used for sampling the spin configurations in HAN 3D – draw_sample. With some small modifications (e.g. dimension of tensors), it can be used for sampling spins in any number of dimensions. The function follows the points 1-3. from section 3.1.
The draw_sample function calls several functions that are implemented explicitly for three dimensions. For example, find_borders (Listing 3) takes as input a cube of spins and returns the values of spins on its walls (but excluding the edges). The function make_cross (Listing 4) takes as input the spin values and forms from them the 3D cross-like structure (see green spins in Figure 1) - the output of the function is a cube with zeros everywhere except for the cross-like structure where $\pm 1$ values are substituted. The function divide_into_cubes (Listing 5) takes as an input a cube and returns a list of eight cubes with edges twice shorter than the original ones. There are several other functions used in the algorithm that need to be implemented for each dimension separately, unless some general versions of these functions are developed.
def draw_sample(
net_b, int_nets,
beta: float, L: int, batch_size: int,
device: torch.device, float_dtype=torch.float32):
# Sampling the three faces of the original lattice
sample_b = net_b.sample(batch_size)
list_args_for_nets: list[list[torch.Tensor]] = [sample_b]
sample: torch.Tensor = create_cube(sample_b, L, L, L, device)
# sample is now a (L+1)^3 cube containing the faces of the original lattice
# taken from sample_b, with opposite faces being duplicates of each other.
sample, list_args_for_nets = han(sample,
int_nets, list_args_for_nets, device)
# heat bath
border = find_borders(sample)
corona = torch.sum(border, dim=(2, 3))
spins, log_prob_heatbath = heatbath(corona, beta, float_dtype)
sample[:, 0, 1, 1, 1] = spins[:, 0]
# recreation of original cubic sized sample
# from its subdivision into small cubes.
log_prob_heatbath = log_prob_heatbath.view(-1, batch_size).sum(0)
while sample.shape[0] > batch_size:
sample = add_into_cube(
torch.chunk(sample, 8, dim=0))
# delete the last faces, as they are duplicates of the first.
sample = sample[:, :, :-1, :-1, :-1]
return sample, list_args_for_nets, log_prob_heatbath
Listing 1: The main function used to sample configurations with HAN 3D. It creates a batch of 3D configurations of $L^{3}$ size each. Apart from the configurations, it returns the log of the probability, but only of the spins generated using the heat bath. This, together with a list of inputs to all the networks, enables us to compute the probability of the whole configuration. The heat bath generation on the last level of the hierarchy is done by the heatbath function. The hierarchical generation is handled by the han function described later. This function splits its input into small ($3^{3}$) cubes that have to be reassembled later by successive applications of the add_into_cube function.
def han(sample, nets, list_args_for_nets, device) -> tuple[torch.Tensor, list[list[torch.Tensor]]]:
b_size = sample.shape[0]
l = sample.shape[2] - 1
for net in nets:
l //= 2
b_size *= 8
border = find_borders(sample)
cross = net.sample(border)
list_args_for_nets.append([cross, border])
sample += make_cross(
cross, sample.shape[2], sample.shape[3], sample.shape[4], device)
sample = torch.cat(divide_into_cubes(sample), dim=0)
return sample, list_args_for_nets
[…]
def find_borders(cube: torch.Tensor) -> torch.Tensor:
assert (cube.shape[2]==cube.shape[3]==cube.shape[4])
t = cube.shape[2] - 1
borders = torch.cat(
[cube[:, :, 0, 1:t, 1:t],
cube[:, :, 1:t, t, 1:t],
cube[:, :, t, 1:t, 1:t],
cube[:, :, 1:t, 0, 1:t],
cube[:, :, 1:t, 1:t, 0],
cube[:, :, 1:t, 1:t, t]],
dim=2)
return borders
Listing 3: The function returns a tensor consisting of the boundary spins of given cube. The edges of the cube are discarded, as we are interested only in the spins that have nearest neighbours in the interior of the cube.
def make_cross(cross: torch.Tensor, L: int, device: torch.device):
assert is_odd(L)
cube = torch.zeros(
[cross.shape[0], 1, L, L, L], device=device, dtype=torch.float32
arm_len = L - 2
arm_len2 = (arm_len - 1) // 2
xz_size = arm_len * arm_len;
yz_half_size = arm_len2 * arm_len
yx_quater_size = arm_len2 * arm_len2
cross_parts = torch.split(
cross,
xz_size,
yz_half_size, yz_half_size,
yx_quater_size, yx_quater_size, yx_quater_size, yx_quater_size
dim=2,
# Loading XZ square (L, L)
i = 0
cube[:, :, 1: arm_len + 1, L // 2, 1: arm_len + 1] = cross_parts[i].reshape(
cross.shape[0], 1, arm_len, arm_len)
i += 1
# Loading the two YZ rectangles (L, (L-1)/2)
for y in range(2):
offy = y * (arm_len2 + 1)
cube[:, :, L // 2, 1 + offy: arm_len2 + 1 + offy, 1: arm_len + 1] = (
cross_parts[i].reshape(cross.shape[0], 1, arm_len2, arm_len))
i += 1
# Loading the four XY squares ((L-1)/2, (L-1)/2)
for x, y in product(range(2), range(2)):
offx = x * (arm_len2 + 1)
offy = y * (arm_len2 + 1)
cube[:, :,
offx + 1: offx + arm_len2 + 1, offy + 1: offy + arm_len2 + 1, L // 2] = (
cross_parts[i].reshape(cross.shape[0], 1, arm_len2, arm_len2))
i += 1
return cube
Listing 4: The function takes as input a flat tensor of spins without duplicates, constructs a 3D cross-like structure, and embeds it into an empty (filled with zeros) cube.
def divide_into_cubes(cube: torch.Tensor) -> list[torch.Tensor]:
assert cube.shape[2] == cube.shape[3] == cube.shape[4]
cube_size = cube.shape[2]
small_cube_size = (cube_size + 1) // 2
begin = (0, small_cube_size - 1)
end = (small_cube_size, cube_size)
return [cube[:, :, begin[x]:end[x], begin[y]:end[y], begin[z]:end[z]]
for z in range(2) for x in range(2) for y in range(2)]
Listing 5: The function divides a cube into eights smaller cubes with the overlaping borders.
def add_into_cube(eight_cubes: Sequence[torch.Tensor]):
small_cube_size = eight_cubes[0].shape[2]
cube_size = 2 * small_cube_size - 1
batch_size = eight_cubes[0].shape[0]
cube = torch.zeros(batch_size, 1, cube_size, cube_size, cube_size, device=eight_cubes[0].device)
begin = (0, small_cube_size - 1)
end = (small_cube_size, cube_size)
i = 0
for z in range(2):
for x in range(2):
for y in range(2):
cube[:, :, begin[x]:end[x], begin[y]:end[y], begin[z]:end[z]] = eight_cubes[i]
i += 1
return cube
Listing 6: The function peforming the inverse of the divide_into_cubes function: it takes eight cubes and puts them back into one large cube.